OpenStudy (anonymous):

subtract and write the answer in simpleast form: (9x)/(3x-4)-(12)/(3x-4)

OpenStudy (anonymous):

(9x - 12) / (3x - 4) = 3(3x - 4) / (3x - 4) = 3

OpenStudy (anonymous):

What rule do I use to do this kind of subtraction problem? \[3/5 - 1/5\] It's the same rule for the problem you're working on, since the denominators are the same. Once you've used that rule, factor and cancel terms.